Solve the equation (2/3 (11x - 1/6) - 3/4 = - 1/4)
Answer by themathtutor2009(81)

2/3 (11x - 1/6) - 3/4 = - 1/4


8(11x - 1/6) - 9 = - 3....mult through by 12

88x - 8/6 - 9 = - 3


528x - 8 - 54 = - 18....mult through by 6

528x-62=-18

528x=44

x=44/528

x=1/12